OpenStudy (anonymous):

The sum of five consecutive odd numbers and four consecutive even numbers together is 325. Also, the largest odd number is 9 less than smallest even number. What is the sum of the second largest odd number and the second smallest even number ?

OpenStudy (campbell_st):

well if you let the 1st odd number be x they are x, x+ 2, x +4 , x+ 6 x+8 the 1st even number by y then the numbers are y, y + 2, y +4, y + 6 there sum to 325 so 5x + 20 + 4y + 12 = 325 or simplifying gives 5x + 4y = 293 you know the largest odd is 9 less than the smallest odd so y - (x + 8) = 9 so y - x = 17 so now you have 2 equations in 2 unknowns that need to be solved simulataneously 5x + 4y = 293 y - x = 17 I'd probably use the substitution method let y = x + 17 so 5x + 4(x + 17) = 293 5x + 4x + 68 = 293 you should be able to solve for x, which will be the 1st odd number. then its just a case of substituting to find the smallest even number... and then all the other odd and even numbers. hope this helps.

OpenStudy (campbell_st):

oops should read.. you know the largest odd is 9 less than the smallest even number

OpenStudy (anonymous):

but the answer is correct

OpenStudy (anonymous):

y ur taking odd no as even no

OpenStudy (anonymous):

x+1,X+3,X+5,X+7 is odd consecutive series

OpenStudy (campbell_st):

nope... if the 1st odd number is x then the next odd is 2 more so x + 2 as an example start at 11... add 2 and you get 13, add2 and you get 15, add 2 and you get 17.. so consecutive odd numbers increase by 2 and the reality is my solution gives the correct answers.